pls. help me solve this prob. tnx.

sam paid 30 dollars for 4 markers and 3 files. A marker cost 1.60 dollars less than a file.
a. find the cost of each marker
b. find the cost of each file

Answer by nerdybill(7384)

sam paid 30 dollars for 4 markers and 3 files. A marker cost 1.60 dollars less than a file.
a. find the cost of each marker
b. find the cost of each file
.
Let f = cost of a file
then
f-1.60 = cost of marker
our equation:
4(f-1.60) + 3f = 30
4f-6.40 + 3f = 30
7f-6.40 = 30
7f = 36.40
f = $5.20 (cost of file)
.
Cost of marker:
f-1.60 = 5.20-1.60 = $3.60
.